# FiDeLiS: Faithful Reasoning in Large Language Models for Knowledge Graph Question Answering

### v2

@@ 5 Conclusion
This paper proposes a retrieval-exploration interactive method specifically designed to enhance intermediate steps of LLM reasoning grounded by KGs. The Path-RAG module and the use of deductive reasoning as a calibration tool effectively guide the reasoning process, leading to more accurate knowledge retrieval and prevention of misleading reasoning chains. Extensive experiments demonstrate that our method, being training-free, not only reduces computational costs but also offers superior generality, consistently outperforming established strong baselines across three distinct benchmarks. We believe this study will significantly benefit the integration of LLMs and KGs, or serve as an auxiliary tool to enhance the interpretability and factual reliability of LLM outputs.
Appendix A Definitions
A.1 Knowledge Graph-based Question Answering (KGQA)
In this work, we mainly focus on the question answering task based on the knowledge graph.
Knowledge Graphs (KGs) contain abundant factual knowledge in the form of a set of triples: $\mathcal{G}=$$\left\{\left(e,r,e^{\prime}\right)\mid e,e^{\prime}\in\mathcal{E},r\in\mathcal{R}\right\}$, where $\mathcal{E}$ and $\mathcal{R}$ denote the set of entities and relations, respectively.
Knowledge Graph Question Answering (KGQA) is a typical reasoning task based on KGs. Given a natural language question $q$ and a $\mathrm{KG}$ $\mathcal{G}$, the task aims to design a function $f$ to predict answers $a\in\mathcal{A}_{q}$ based on knowledge from $\mathcal{G}$, i.e., $a=f(q,\mathcal{G})$. Following previous works Luo et al. (2024; 2023), we assume the entities $e_{q}\in\mathcal{T}_{q}$ mentioned in $q$ and answers $a\in\mathcal{A}_{q}$ are labeled and linked to the corresponding entities in $\mathcal{G}$, i.e., $\mathcal{T}_{q},\mathcal{A}_{q}\subseteq\mathcal{E}$.
Appendix B Potential Impacts and Limitations
The proposed method holds the potential to significantly enhance the performance of large language models (LLMs) by tackling the issue of hallucinations, thereby fostering deep, responsible reasoning. By integrating KGs with LLMs, the approach not only facilitates more accurate knowledge retrieval but also leverages deductive reasoning capabilities to steer the reasoning process and circumvent logical fallacies. The method is characterized by its stepwise, generalizable approach, and the use of deductive verification as a stopping criterion, which together may reduce superfluous computation and curb misleading reasoning chains. Moreover, due to its training-free nature and lower computational demands, this method could seamlessly serve as a plug-in for other existing frameworks.
However, it’s important to recognize certain limitations. Since the method is still in its early stages of development, it might encounter unanticipated challenges or drawbacks in real-world applications. Its dependency on external KGs means that the quality and comprehensiveness of these resources can affect its overall effectiveness. Computational constraints might also arise when dealing with very large or intricate graphs. Furthermore, although the method has demonstrated promise in benchmark tests, its performance across more diverse or specialized tasks remains untested.
Appendix C Bottleneck of Beam Search Efficiency
The bottleneck of computation is the beam search process, which contributes to $N*D$ times LLM calling, where $D$ is the depth (or equivalently length) of the reasoning path, and $N$ is the width of the beam-search (how many paths are remained in the pool in each iteration). Specifically, we need to call $ND+D+C$ times LLM for each sample question, where $C$ is a constant (equals to $1$ if there is no error occurs when calling the API). Sun et al. (2023) illustrates that the computational efficiency can be alleviated by replacing LLMs with small models such as BM25 and Sentence-BERT for the beam search decision since the small models are much faster than LLM calling. In this way, we can reduce the number of LLM calling from $ND+D+C$ to $D+C$. However, Sun et al. (2023) shows that this optimization sacrifices the accuracy due to the weaker scoring model in decision making.
Appendix D Experiment Details
D.1 Baselines
Baselines. (1) RoG (Luo et al., 2024): embed structure knowledge graph from KGs into the underlying neural networks during the pretraining and fine-tuning process to generate the reasoning path and explanation. (2) ToG (Sun et al., 2023): ask LLM to iteratively explore multiple possible reasoning paths on KGs until the LLM determines that the question can be answered based on the current reasoning paths. Our method FiDeLiS follows the similar paradigm. (3) NSM (He et al., 2021) utilizes the sequential model to mimic the multi-hop reasoning process. (4) KD-CoT (Wang et al., 2023a) retrieves relevant knowledge from KGs to generate faithful reasoning paths for LLMs. (5) DeCAF (Yu et al., 2022) combines semantic parsing and LLMs reasoning to jointly generate answers, which also reach salient performance on KGQA tasks.
Implementation Details. We set the default beam width as $4$ and depth as $4$ without specific annotation. We set the $\alpha$ in Eq 3 as $0.3$ to ensure reproducibility. For LLMs usage, we set all the inference using temperature $T=0.3$ and $p=1.0$. The hyperparameter tuning experiments for beam search and $\alpha$ can be found in Figure 2 and Table 9.
Backboned LLMs. We assess our approach on closed- and open-source LLMs. For closed-source LLMs, we choose GPT-4-turbo, GPT-3.5-turbo [^1] We use the recent gpt-4-turbo model that is released in 2024-04-09 from https://platform.openai.com/docs/models/gpt-4-turbo-and-gpt-4 and gpt-3.5-turbo model notated as “gpt-3.5-turbo-0125” from https://platform.openai.com/docs/models/gpt-3-5-turbo. to report and compare the results on all datasets. We use Llama-2-13B (Touvron et al., 2023) and Mistral-7B (Jiang et al., 2023a) as our open-source LLMs to conduct cost–performance analysis on different datasets. The experiment results of open-source models can be found in Table 10.
Dataset #Ans = 1 2 $\geq$ #Ans $\leq$ 4 5 $\geq$ #Ans $\leq$ 9 #Ans $\geq$ 10
WebQSP 51.2% 27.4% 8.3% 12.1%
CWQ 70.6% 19.4% 6% 4%
D.2 Datasets
@@
Datasets & Metrics. We adopt three benchmark KGQA datasets: WebQuestionSP (WebQSP) (Yih et al., 2016), Complex WebQuestions (CWQ) (Talmor & Berant, 2018) and CR-LT-KGQA (Guo et al., 2024) in this work. We follow previous work (Luo et al., 2024) to use the same training and testing splits for fair comparison over WebQSP and CWQ. The questins from both WebQSP and CWQ can be reasoned using Freebase KGs [^2] https://github.com/microsoft/FastRDFStore. To address the bias in WebQSP and CWQ, which predominantly feature popular entities and there is a likelihood that their data might have been incorporated into the pre-training corpora of LLMs, we further test our method on CR-LT-KGQA (further discussed in Appendix Section 8). We use the complete dataset from CR-LT-KGQA in our experiments, as it comprises only 200 samples. Each of the question can be reasoned based on the Wikidata [^3] https://www.wikidata.org/wiki/Wikidata:Main_Page. The statistics of the datasets are given in Table 7 and Table 8. To streamline the KGs, we utilize a subgraph of Freebase by extracting all triples that fall within the maximum reasoning hops from the question entities in WebQSP and CWQ followsing RoG (Luo et al., 2024). Similarly, we construct the corresponding sub-graphs of Wikidata for CR-LT-KGQA as well. We assess the performance of the methods by analyzing the F1 and Hits@1 metrics for CWQ and WebQSP, and by evaluating the accuracy for CR-LT-KGQA.
@@
Motivation of CR-LT-KGQA. The motivation for evaluating over CR-LT-KGQA is that the majority of existing KGQA datasets, including WebQSP and CWQ, predominantly feature popular entities. These entities are well-represented in the training corpora of LLMs, allowing to often generate correct answers based on their internal knowledge, potentially without external KGs. Moreover, since WebQSP and CWQ have been available for several years, there is a likelihood that their data might have been incorporated into the pre-training corpora of LLMs, further reducing the need for external KGs during question-answering.
@@
Against this backdrop, we specifically chose CR-LT-KGQA for our evaluation as it includes queries related to obscure or long-tail entities, the distribution of the frequency and entity popularity of CR-LT is shown as Figure 4. Such scenarios are where KGs play a critical role because they provide a reliable source of verifiable information, crucial when LLMs encounter entities not well covered in their training data. By testing our methods on CR-LT-KGQA, we aim to explore how effective LLMs can operate when combined with KGs, especially in contexts involving less common knowledge domains where LLM performance typically declines. This evaluation helps us understand the extent to which KGs remain necessary for supporting LLMs in a diverse range of query scenarios.
D.3 Parameter Tuning for $\alpha$ for Scoring Function
For hyperparameter tuning regarding $\alpha$ for Eq 3, we added an extra comparison in Table 9 which shows that $\alpha$ is actually not a very impact parameter for the whole system, however for the reproducibility of our method, we set the $\alpha$ as 0.3.
@@ D.3 Parameter Tuning for $\alpha$ for Scoring Function
Table 9: Parameter tuning for $\alpha$ for scoring function over WebQSP
$\alpha$ 0.1 0.2 0.3 0.4 0.5 0.6 0.7 0.8 0.9 1.0
Hits@1 83.76 84.15 84.39 83.98 83.55 84.12 83.79 82.25 83.60 83.73
Appendix E Algorithm Description
@@
Appendix F Prompt List
F.1 Plan-and-solve
In-Context Few-shot
Q: {Query}
A:
F.2 Deductive-verification
In-Context Few-shot
A:
F.3 Adequacy-verification
In-Context Few-shot
A:
F.4 Beam Search
In-Context Few-shot
A:
F.5 Reasoning
In-Context Few-shot
Question: {Query}
Reasoning path: {reasoning_path}
Only return the answer to the question.
A:
F.6 Demonstration of Deductive Verification
Table 10: Comparison of FiDeLiS using different backbone LLMs.
Backend Models Methods WebQSP CWQ CR-LT
Hits@1 (%) F1 (%) Hits@1 (%) F1 (%) Acc (%)
Llama-2-13B FiDeLiS 72.34 69.78 58.41 54.78 60.87
Mistral-7B FiDeLiS 74.11 70.23 60.71 56.87 63.12
Appendix G Related Works: Reasoning with LLM Prompting
With the development of LLMs, many creative ways to leverage LLMs are proposed: End-to-End, Chain-of-Thought, and Semantic Parsing/Code Generation.
End-to-End methods (Chen, 2022; Hoffmann et al., 2022) aims to leverage LLMs to generate final answers directly, often done by providing a task description and/or a few examples for in-context learning. While it offers convenience, however, it suffers from the un-interpretability of the generation, and the lack of explicit steps and reliance solely on the LLMs’ training data may result in un-robustness and exhibit sensitivity to slight input variations.
Chain-of-Thought methods (Wei et al., 2022a; Kojima et al., 2022; Gong et al., 2020) emphasize breaking down complex reasoning into a series of intermediate steps to support complex reasoning. However, CoT suffers from unreliability and uncontrollability since the generated reasoning steps may not always align with the intended logical progression or may produce incorrect or inconsistent answers when multiple questions are posed consecutively.
Plan-and-solve methods (Wang et al., 2023b) prompts LLMs to generate a plan and conduct reasoning based on it. DecomP He et al. (2021) prompts LLMs to decompose the reasoning task into a series of sub-tasks and solve them step by step.
Semantic Parsing/Code Generation methods (Cheng et al., 2023; Ye et al., 2023; Gemmell & Dalton, 2023) leverages LLMs to convert natural language queries into executable code or structured representations. This approach enables more precise control over the output and facilitates better interpretability. However, it has its limitations, such as the limited coverage of programming language grammar and semantics. This restricts the model’s ability to handle complex programming tasks and may require additional techniques to handle out-of-domain queries effectively. Cheng et al. (2023) propose to leverage LLMs to bridge the out-of-domain queries, however, the form of knowledge injection still lacks a thorough exploration and the trigger functions are still naive for complex queries. However, the problem of hallucinations and lack of knowledge affect the faithfulness of LLMs’ reasoning. ReACT Yao et al. (2022) treats LLMs as agents, which interact with the environment to get the latest knowledge for reasoning. To explore faithful reasoning, FAME Hong et al. (2023) introduces the Monte-Carlo planning to generate faithful reasoning steps. RR He et al. (2022) and KD-CoT Wang et al. (2023a) further retrieve relevant knowledge from KGs to produce faithful reasoning plans for LLMs.

### v3

@@ 6 Conclusion
This paper proposes a retrieval-exploration interactive method specifically designed to enhance intermediate steps of LLM reasoning grounded by KGs. The Path-RAG module and the use of deductive reasoning as a calibration tool effectively guide the reasoning process, leading to more accurate knowledge retrieval and prevention of misleading reasoning chains. Extensive experiments demonstrate that our method, being training-free, not only reduces computational costs but also offers superior generality. We believe this study will significantly benefit the integration of LLMs and KGs, or serve as an auxiliary tool to enhance the interpretability and factual reliability of LLM outputs.
Limitations
Our work demonstrates a promising advancement by integrating KGs with LLMs to reduce hallucinations and promote deep, faithful reasoning through deductive verification. However, the method exhibits certain limitations. Its reliance on external KGs means that the overall effectiveness is contingent on the quality and comprehensiveness of these resources, and challenges may arise when encountering incomplete, inconsistent or outdated information. Despite these limitations, the open-KGs like Wikidata and DBpedia used in our study are of high quality, benefiting from years of updates by an extensive community. For domain-specific KGs, although there may currently be gaps in quality, we are optimistic about future enhancements. Given the significant societal impact and the noticeable boosts in LLM performance facilitated by KGs, it is likely that community efforts will continue to refine and expand these resources.
Appendix A Experiment Details
A.1 Baselines
We consider the following methods including training-free (highlighted with *) and training-based methods as baselines:
  • NSM He et al. (2021) propose a teacher-student approach for KGQA task, where the student network aims to find the correct answer to the query, while the teacher network tries to learn intermediate supervision signals for improving the reasoning capacity of the student network.
  • KD-CoT Wang et al. (2023) propose to verify the sub-reasoning process of LLMs through the external KGs to facilitate faithful reasoning.
  • DeCAF Yu et al. (2023) use a text-based retrieval instead of entity linking to select question-related information from the KG, and generate logical forms and direct answers respectively. They combine the logical-form-executed answers and directly-generated answers to obtain the final output.
  • KAPING Baek et al. (2023) proposes a zero-shot knowledge-augmented prompting method. It first retrieves triples related to the question from the graph, then prepends them to the input question in the form of a prompt, which is then forwarded to LLMs to generate the answer.
  • ToG Sun et al. (2023): conduct the reasoning on KGs by iteratively exploring multiple potential reasoning paths and concludes the final answer by aggregating the evidence from retrieved reasoning paths.
  • RoG Luo et al. (2024a): incorporate the underling structure of KGs into LLMs throught pre-training and fine-tuning to generate the reasoning path and explanation.
  • GCR Luo et al. (2024b) propose to integrate KG structure into the LLM decoding process to conduct graph-constrained reasoning.
A.2 Datasets & Metrics
@@
We consider three KGQA benchmark: WebQuestionSP (WebQSP) Yih et al. (2016), Complex WebQuestions (CWQ) Talmor and Berant (2018) and CR-LT-KGQA Guo et al. (2024) in this work. We follow previous work Luo et al. (2024a) to use the same training and testing splits for fair comparison over WebQSP and CWQ. The questions from both WebQSP and CWQ can be reasoned using Freebase KGs [^1] https://github.com/microsoft/FastRDFStore. To address the bias in WebQSP and CWQ, which predominantly feature popular entities and there is a likelihood that their data might have been incorporated into the pre-training corpora of LLMs, we further test our method on CR-LT-KGQA (discussed in §A.2). We use the complete dataset from CR-LT-KGQA in our experiments, as it comprises only 200 samples. Each of the question can be reasoned based on the Wikidata [^2] https://www.wikidata.org/wiki/Wikidata:Main_Page. The statistics of the datasets are given in Table 10 and Table 9. To streamline the KGs, we follow RoG Luo et al. (2024a) and utilize a subgraph of Freebase by extracting all triples that fall within the maximum reasoning hops from the question entities in WebQSP and CWQ. Similarly, we construct the corresponding sub-graphs of Wikidata for CR-LT-KGQA. We assess the performance of the methods by analyzing the F1 and Hits@1 metrics for CWQ and WebQSP, and by evaluating the accuracy for CR-LT-KGQA. The statistics of the datasets can be found in Table 9 and Table 10.
@@
Dataset #Ans = 1 2 $\geq$ #Ans $\leq$ 4 5 $\geq$ #Ans $\leq$ 9 #Ans $\geq$ 10
WebQSP 51.2% 27.4% 8.3% 12.1%
CWQ 70.6% 19.4% 6% 4%
Motivation of CR-LT-KGQA.
The motivation for evaluating over CR-LT-KGQA is that the majority of existing KGQA datasets, including WebQSP and CWQ, predominantly feature popular entities. These entities are well-represented in the training corpora of LLMs, allowing to often generate correct answers based on their internal knowledge, potentially without external KGs. Moreover, since WebQSP and CWQ have been available for several years, there is a likelihood that their data might have been incorporated into the pre-training corpora of LLMs, further reducing the need for external KGs during question-answering. To this end, we utilize the CR-LT-KGQA benchmark, which features queries specifically crafted to target obscure and long-tail entities. Figure 4 illustrates the distribution of entity frequency and popularity in CR-LT, underscoring the inherent challenges of these queries. In such scenarios, knowledge graphs are indispensable as they offer a reliable, verifiable source of information, particularly for entities that are poorly represented in the training data of large language models. By testing our methods on CR-LT-KGQA, we investigate the extent to which integrating KGs can bolster LLM performance in less common knowledge domains, where their effectiveness typically declines. This evaluation not only demonstrates the potential synergy between LLMs and KGs but also clarifies the critical role that KGs continue to play in supporting LLMs across diverse query scenarios.
@@
Figure 5: Parameter tuning for $\alpha$ for scoring function over WebQSP
Backend Models WebQSP CWQ CR-LT
Hits@1 (%) F1 (%) Hits@1 (%) F1 (%) Acc (%)
Llama-2-13B 72.34 69.78 58.41 54.78 60.87
Mistral-7B 74.11 70.23 60.71 56.87 63.12
Table 11: Performance over Open-sourced LLMs.
A.3 Backbone LLMs & Embedding Methods
Backbone LLMs. We assess our approach on closed- and open-source LLMs. We consider closed-source models like GPT-4-turbo (between Feb, 2024 to July, 2024), GPT-3.5-turbo (between Feb, 2024 to July, 2024), GPT-4o, GPT-4o-mini (between Nov, 2024 to Jan 2025) from OpenAI, and open-sourced models like meta-llama-2-13B from Meta and mixtral-7B from Mixtral AI. The experiment results of open-source models can be found in Table 11. We set all the inference configs using temperature $T=0.3$ and $p=1.0$.
Embedding Methods. We assess the robustness of the retrieval module Path-RAG on different embedding models. We consider probabilistic ranking function like BM25333https://en.wikipedia.org/wiki/Okapi_BM25, dense retrieval using smaller language models like SentenceBERT [^4] https://sbert.net and E5555https://huggingface.co/intfloat/e5-large, and more advanced embedding model like text-embedding-3-small from OpenAI [^6] https://platform.openai.com/docs/guides/embeddings.
A.4 Implementation Details
We set the default beam width as $4$ and depth as $4$ without specific annotation. We set the $\alpha$ in Eq 2 as $0.3$ to ensure reproducibility. For hyper-parameter tuning regarding $\alpha$ for Eq 2 and beam search width and length, we conduct experiments as shown in Figure 5 and Figure 6.
Analysis of Beam Search.
We investigate the effect of hyper-parameters like beam width and depth in beam search, as illustrated in Figure 6. By varying the width and depth from 1 to 4, we observe that overall performance improves as both parameters increase, peaking when the search depth exceeds 3 for the WebQSP and CWQ datasets. However, beyond a depth of 3, performance begins to decline, likely because only a small fraction of questions in these datasets require reasoning at greater depths. In contrast, increasing the beam width consistently enhances performance, highlighting the benefits of broader exploration in search.
(a) RD over CWQ
(b) RD over WebQSP
(c) BW over CWQ
(d) BW over WebQSP
Figure 6: Analysis of various beam search width (BW) and reasoning depth (RD).
A.5 Robustness Analysis Across Different Domains and KGs
KGs vary in structure and domain-specific characteristics, so consistent performance across both general and specialized KGs can reflect a method’s adaptability to diverse real-world applications. To this end, we conduct robustness analysis of FiDeLiS across different domains and KGs to verify the generalizability. To perform this analysis, we introduced a new dataset, MedQA-USMIE, sourced from MedQA Jin et al. (2020), which is designed to require domain-specific biomedical and clinical knowledge. The dataset is a 4-way multiple-choice question-answering task, and we extracted 300 examples from its test set for evaluation. The corresponding biomedical KG is based on Disease Database and DrugBank Zhang et al. (2022). The results, presented in Table 2, indicate that our method exhibits consistent robustness across different types of KGs. Our scoring function, enhanced by incorporating next-hop neighbor information $S_{r}^{i}+S_{e}^{i}+\alpha\max_{\forall j\in N_{i}}(S_{r}^{j}+S_{e}^{j})$, achieves higher performance gains in both WebQSP and MedQA-USMIE, particularly improving accuracy in the specialized biomedical domain. These findings validate that our method can effectively handle the challenges posed by both general and domain-specific knowledge graphs, indicating strong adaptability and robustness.
Method WebQSP MedQA-USMIE
ToG 81.84 42.37
[…]
Path-RAG w/ $S_{r}^{i}+S_{e}^{i}$ 83.15 44.31
Path-RAG w/ $S_{r}^{i}+S_{e}^{i}+\alpha\max_{\forall j\in N_{i}}(S_{r}^{j}+S_{e}^{j})$ 84.39 46.45
Table 12: Robustness analysis of our method across different domains
Appendix B Bottleneck of Beam Search Efficiency
The bottleneck of computation is the beam search process, which contributes to $N*D$ times LLM calling, where $D$ is the depth (or equivalently length) of the reasoning path, and $N$ is the width of the beam-search (how many paths are remained in the pool in each iteration). Specifically, we need to call $ND+D+C$ times LLM for each sample question, where $C$ is a constant (equals to $1$ if there is no error occurs when calling the API). Sun et al. (2023) indicate that the computational efficiency can be alleviated by replacing LLMs with small models such as BM25 and Sentence-BERT for the beam search decision since the small models are much faster than LLM calling. In this way, we can reduce the number of LLM calling from $ND+D+C$ to $D+C$. However, this may sacrifices the accuracy due to the weaker scoring model in decision making Sun et al. (2023).
Appendix C Prompt List
C.1 Plan-and-solve
In-Context Few-shot
Q: {Query}
A:
C.2 Deductive-verification
In-Context Few-shot
A:
C.3 Adequacy-verification
In-Context Few-shot
A:
C.4 Beam Search
In-Context Few-shot
A:
C.5 Reasoning
In-Context Few-shot
Question: {Query}
Reasoning path: {reasoning_path}
Only return the answer to the question.
A:
C.6 Demonstration of Deductive Verification